OpenStudy (anonymous):

I am having trouble steps to solving this problem (need step by step) x+12 4 ------ = ----- x 3

OpenStudy (anikate):

first multiply 3x to both sides, then tell me what u get?

OpenStudy (anonymous):

3(X+12)+3X+36 3X(4)=12 ?

OpenStudy (anikate):

no let me show you

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

3x+36=4x therefore x=36

OpenStudy (anonymous):

x=36

OpenStudy (anonymous):

what happen to 3x and 4x? I had the answer just didn't know the steps to figure it out

OpenStudy (anikate):

x+12 4 ------ = ----- x 3 3(x+12)=4x 3x+36=4x 36=4x-3x x=36

OpenStudy (anikate):

those r all the steps

OpenStudy (anonymous):

ok thank you ever so much (haven't done math in 23 years and just recently tested to go into RN program but needed Algebra Class

OpenStudy (anikate):

oh ok.. well yw. and may I plz get a medal :D

OpenStudy (anonymous):

what do I do to give you one of those? First time on here

OpenStudy (anikate):

Oh ok right across from my name on this chatbox you will see a medal just click it

OpenStudy (anikate):

oh thx

OpenStudy (anikate):

I gave u a medal for good question when u posted it

OpenStudy (anonymous):

Thanks for your help

OpenStudy (anikate):

YW GOOD LUCK FOR YOUR TEST!!!!!! :D BYE cya :)